Given that the price in a competitive industry is 100 the firm's total cost equation is given as TC=50+20q+q2
Determine the firm's stock total re




Given that the price in a competitive industry is 100 the firm's total equation is given as TC=50+20q+q2
Determine the firm's total stock revenue
The profit equation, the quantity which maximizes the profit , profit that maximizes the quantity what is the total cost the total revenue and the profit
1
Expert's answer
2021-08-09T15:52:48-0400

"TC=50+20q+q^2"

"P = 100"

"Profit = Revenue - Total \space cost"

"Profit = 100q - (50+20q+q^2)"

"Revenue = Price\times Quantity"

"Revenue = 100\times q"

"MC = 20 +2q"

At profit max.

"Price = MC"

"100 = 20 +2q"

"80 = 2q\\

q = 40"


"Revenue = 100\times40"

"Revenue = 4000"


Profit max. quantity "40" and revenue "4000"